# Function-based Labels for Complementary Recommendation: Definition, Annotation, and LLM-as-a-Judge

### v1

@@ 5.3 Classification Accuracy of LLMs
    LR: Reference performance in a ML setting.
5.4 Comparison between Human and LLMs FBLs
Figures 2 and 3 show the confusion matrices comparing human-annotated FBLs with those generated by Llama and GPT in the 9-cls setting. In Llama, outputs for complementary relationships are concentrated on B-1 (“Item $x$ can be replenished with item $y$”) indicating a relatively good classification of complementary relationships, but it struggles with distinguishing more complex relationships and their directions, such as replenishment and combination. In contrast, GPT shows a reduced tendency for localization in complementary relationships, suggesting that it can capture more complex item relationships than Llama. Both LLMs also tend to classify item pairs that humans labeled as “A. Items $x$ and $y$ have the same function and usage.” and “C-4. Combining $x$ and $y$ makes them more useful.” as “D. Items x and y have no relationship.” Examples of the former include pairs, such as envelopes of different sizes or pens with different ink properties, including mechanical pencils and oil-based pens, which are difficult for humans to judge as substitutes. Examples of the latter include pairs such as sponges and bleach or disposable chopsticks and soup cups, which can be used together in the same scene (kitchen or dining), but do not necessarily have to be used, indicating a weak relationship. The analysis of these item pair trends suggests that LLMs tend to diverge from human judgment for item pairs with ambiguous relationships or differences, functioning as stricter annotators than humans by classifying such pairs as unrelated.
Figure 2: Correspondence between Llama-generated and human-annotated labels (maximum value set to 100)
Figure 3: Correspondence between GPT-generated and human-annotated labels (maximum value set to 100)
6 Conclusion
In this study, we propose FBLs that focus on the functional relationships between items as a new definition of complementary relationships independent of user behavior history. We constructed a human-annotated dataset based on the FBLs relationship definitions and evaluated the classification accuracy of these relationships using ML models and LLMs. The label distribution results of the human-annotated FBL dataset clarified the existing relationships in complementary recommendations from a functional perspective. Additionally, labeling by LLMs based on FBLs improved the response consistency, suggesting that FBLs function as a clear definition of complementary relationships (RQ1). In the evaluation experiment of ML models using FBLs, LR and SVM achieved an average macro-F1 of approximately 0.82 in both Valid and Test, demonstrating high classification accuracy (RQ2). In the evaluation experiment of relationship estimation by LLMs, FBLs assigned by LLMs achieved a high macro-F1 of up to 0.849 compared with human-annotated FBLs, suggesting that LLMs can function effectively as annotators mimicking human perception when using FBLs (RQ3).
However, concluding that the FBLs defined in this study cover all possible relationships, and more comprehensive and clear relationship definitions are required. In addition, although this study confirmed the usefulness of LLMs as annotators, applying LLMs to a vast number of item pairs incurs high costs in terms of time and money. Therefore, to reduce these costs, the development of automatic pair classification using ML models is required. In particular, active learning methods that utilize teacher labels obtained from LLMs [36, 37] are considered promising approaches and directions for future research.
Acknowledgments

### v2

@@ 6 Conclusion
    The current dataset is limited in size and category coverage; larger-scale annotation with additional annotators and diverse item domains will be necessary to confirm reproducibility and generalizability.
  • 2.
[…]
  • 3.
    The present study was conducted in Japanese; extending the dataset and experiments to other languages will allow us to consider cultural and linguistic influences.
  • 4.
    Expanding the dataset to include long-tail items will enable examination of domain shifts that naturally occur in real-world e-commerce.
Although this study confirmed the usefulness of LLMs as annotators, applying LLMs to a vast number of item pairs incurs high costs in terms of time and money. Therefore, to reduce these costs, the development of automatic pair classification using ML models is required. In particular, active learning methods that utilize teacher labels obtained from LLMs [28, 29] are considered promising approaches and directions for future research. Building on this perspective, it is also important to recognize that LLMs may not always possess sufficient knowledge to provide highly reliable labels. Therefore, exploring active learning paradigms that explicitly account for uncertain labeling knowledge [30] could further enable more robust and trustworthy utilization of LLM-generated FBLs.
Funding
@@
Appendix A LLM Prompts for 9-class (English Version)
The prompt used for the 9-class annotation is shown below.
Input: You are a highly capable assistant that assigns a unique label to the relationship between #ItemA and #ItemB by strictly following the #Procedure. In your output, do not include anything other than the label information in the format "{’label’: <corresponding option number>}".
# Procedure
(1) Based on the detailed information of #Item A and #Item B, understand how each item is used.
(2) Go through the options in the #Flowchart (1) to (9) in order, and decide which option best matches the relationship between #ItemA and #ItemB.
(3) Output the corresponding option number strictly in the JSON format defined in #Output Format. Do not add any explanation or extra text.
# Flowchart
(1) Items A and B have the same function and usage.
(2) Item A can be replenished with item B.
(3) Item B can be replenished with item A.
(4) Item A and By must be combined to be usable.
(5) When combined with item B, item A becomes more useful.
(6) When combined with item A, item B becomes more useful.
(7) Combining A and B makes them more useful.
(8) Items A and B have no relationship.
(9) Items A and B seem to have a relationship, but it is difficult to verbalize.
(10) Others.
# Item A
{Input Item A Descriptions}
# Item B
{Input Item B Descriptions}
# Output Format
{’label’: <corresponding option number>}
Output: